what is (x-1)ln4=2xln3
Answer by stanbon(75887)

what is (x-1)ln4 = 2xln3
---
x*ln(4) - ln(4) = 2x*ln(3)
-------
Get the "x" terms on one side and factor it out to get:
x[ln(4)-2ln(3)] = ln(4)
x[ln[4/9] = ln(4)
----
x = [ln(4)]/[ln(4/9)]
----
x = -1.7095
===========
Cheers,
Stan H.